An 11.7kg block resting on a frictionless, level surface, is pulled with a force of 30N acting at an angle of 30deg above the horizontal axis.
a. Calculate the magnitude of the net force on the block in the x axis.
b. Calculate the magnitude of the net force on the block in the y axis.
c. Calculate the magnitude of the weight of the block.
d. The normal force is _______ the weight of the block.
e. Calculate the magnitude of the Normal Force between the block and the ground.
f. Calculate the magnitude of the horizontal acceleration of the block.
1
Expert's answer
2020-12-14T12:14:50-0500

a. The magnitude of the net force on the block in the x axis is the horizontal projection of "F= 30N" (since it is the only force acting along the x-axis). Thus, obtain:


"F_x = F\cos30\degree \approx 26.0N"

b. The magnitude of the net force on the block in the y axis is zero, since block doesn't move in vertical direction.

c. The magnitude of the weight of the block is:


"P = mg-F\sin 30\degree = 11.7kg\cdot 9.8m/s^2 - 30N\cdot 0.5 = 99.7N"

d. The normal force is equal to the weight of the block.

e. The magnitude of the Normal Force between the block and the ground is:


"N = 99.7N"

f. The magnitude of the horizontal acceleration of the block can be found from the second Newton's law:


"a = \dfrac{F_x}{m} = \dfrac{26.0N}{11.7kg} = 2.2 m/s^2"